# Gravitational reduction of the wave function through the quantum theory of motion

### v1

@@ Abstract
A new perspective based on Bohmian trajectories is provided for the gravitational reduction of the wave function. The quantum motion of both a point particle and an object is investigated in the presence of quantum and gravitational forces, which are related to the probability distribution of the system in its configuration space. We discuss different regimes of motion in this context. By studying the dynamics of a particle or an object, we identify the critical quantities for the transition from the quantum world to the classical world. Furthermore, the reduction time of the wave function is defined using the dynamics of the particle or object. In this method, we aim to express related concepts in an intuitive and simple way.
@@ 1 Introduction
The reduction of the wave function, which is one of the unsolved problems in quantum physics, is studied in different ways. For example GRW theory which proposes that wave function collapse happens spontaneously and randomly over time [1, 2]. Decoherence theory, which describes how interactions with the environment cause a quantum system to lose its coherence and behave in a classical manner[3, 4, 5]. Many-worlds interpretation which suggests that all possible outcomes of a measurement actually occur, each in its own separate branch of the universe. This leads to a vast multiverse where every possible result happens in some branch[6, 7]. The approach we discuss, is the approach of gravitational reduction of the wave function. This approach has been extensively explored in standard quantum mechanics. For a review, see Ref. [8]. But we study it in the framework of Bohmian quantum mechanics as an approach that can provide us with more intuition. Although this may be a matter of taste, the concepts in the Bohmian quantum mechanics allow us to have a clearer picture of what is happening.
The mass of an object has a significant role in determining the boundary between the classical and quantum worlds. In particular, we know that the Planck mass is often considered as a critical threshold. Thus, gravity comes into play. In this interpretation, the self-gravity of the object or the curvature of spacetime caused by its mass directs the object to a particular state. This is not a new word, and it has been about sixty years since the first formulation was presented to explain the problem[9, 10, 11]. For masses higher than the Planck mass ($\sim 10^{-5}g$), the behavior of the particle becomes classical, and for lower masses, it is pushed towards quantum behavior. In this regard, estimates are made for the critical width of the wave packet or the reduction time of the wave function. But, despite these achievements, an ontological and deterministic description of the problem is not provided yet. We cannot predict with certainty to which of the possible states the quantum system will be driven after the measurement. And we resort only to statistical predictions. We are dealing with a concept called the ”wave function” associated with any object and its nature is unknown yet. However, the assumption of such a being has so far helped to make many possible predictions of the quantum world. But, this is not enough to describe the problem in an ontological way. This problem also exists in the Bohmian quantum mechanics, but instead, an attempt has been made to have a deterministic description of the quantum world. In the Bohmian quantum mechanics the role of the wave function is to guide particles on well-defined trajectories. The wave function is affected by the classical potentials and particle mass, but there are no interactions like the interactions between energy fields. In other words, the wave function adjusts itself to matter and the environment, but this does not take place in the form of energy or force exchange between the wave function and matter. At this level, the term ”pilot wave” really makes sense. In addition, the non-local nature of the wave function also increases the complexity of the problem. In practice, the wave function is used as a tool for possible predictions, and under these conditions, it is impossible to present an ontological theory about the reduction of the wave function. None of the interpretations related to the reduction of the wave function can exactly predict the final state of the system after the reduction. Nevertheless, Bohmian quantum mechanics partly helps us to visualize the problem and, due to its structure which involves concepts such as trajectory, velocity and force, we decided to investigate the problem in the Bohmian quantum mechanics.
Figure 1: This figure shows the particle distribution $\rho=|\psi|^{2}$, in configuration space and its quantum gravitational force schematically.
We shall investigate the quantum motion of a particle or an object, when it moves under the influence of quantum force and quantum gravitational force (which will be defined later) in the ensemble of trajectories. The term ”quantum gravitational force” should not be misleading. This expression refers to the force that is defined in the configuration space of a particle where its quantum probability distribution is located. Figure (1) shows the particle distribution in configuration space and quantum gravitational force, schematically.
@@
First, let’s see what gravitational approaches exist in the standard quantum mechanics. According to the uncertainty principle, the wave function of the particle spreads with time and as a result the uncertainty in the particle’s location tends to increase. While gravity tends to localize and reduce the uncertainty in the location of the particle. Basically, the reduction of the uncertainty in the location of the particle by gravity can be related to the concept of reduction of the wave function. In the standard quantum mechanics, gravity enters the problem in such a way that it requires a set of metrics or spacetimes. The superposition of particle states is not stable in the presence of gravity (technically, an ensemble of spacetime metrics) and after a timescale which can be estimated by uncertainty in the particle’s gravitational energy, the superposition breaks down to one of the possible states. The first significant work is the work of Karolyhazi, in which an uncertainty is considered for the metric of spacetime. To achieve the classical regime, the uncertainty in spacetime is minimized in the presence of a scalar field. This leads to obtaining the critical conditions for the transition from the quantum domain to the classical domain. Karolyhazy’s relations for critical width and reduction time of an elementary particle are given by
@@
Here, the reduction time is the time it takes for a quantum state to undergo spontaneous localization, leading to a classical-like behavior. For a proton, one can find $\sigma_{c}\approx 10^{22}m$ and $\tau\approx 10^{53}s$. For an object of mass $M$ and size $R$, Karolyhazy derives a formula as follows:
@@
For a tennis ball of radius $\sim 4cm$ and mass $\sim 57g$, the critical width and reduction time are approximately $\sigma_{c}\sim 10^{-17}cm$ and $\tau\sim 10^{-6}s$ respectively. For details, see[9, 10, 11] .
The gravitational approach of Diosi also determines a critical width for transition from the quantum to the classical world. The first work of Diosi is an investigation through the Schrödinger-Newton equation, which is a nonlinear equation with a self-gravitational potential. It is believed that such a nonlinear term breaks down the unitary transformation and transforms a pure state into a mixture of states. By using the Schrödinger-Newoton equation, the critical width of a stationary wave packet is obtained. Subsequent works by Diosi, deal with the concept of intrinsic quantum imprecision in the spacetime metric. Through the study of fluctuations of the gravitational field when it is measured by a quantum probe over a period of time, the damping time or reduction time of a quantum state in the presence of gravity is defined[12, 13, 14]. The critical width for transition from the quantum world to the classical world of an object, in Diosi’s approach, are as follows:
@@
A macro object refers to a macroscopic body that exhibits classical behavior due to its large mass and size and does not exhibit significant quantum superpositions over macroscopic distances. A micro object exhibits quantum behaviors such as superposition or entanglement. We expect a micro object to be in different places or states at the same time and the width of its wave packet to be greater than its classical size. For a proton, the critical width is about $10^{4}m$, which is really greater than its classical size,.i.e $10^{-15}m$. The damping time or the reduction time for a proton is about $10^{15}s\approx 32000000$ years, which indicates that we can not see the collapse of the proton normally.
The approach of Penrose is very fundamental. In Ref. [15], an effort is made to bring quantum mechanics closer to the principles of general relativity. Penrose shows that in order to establish the principles of general relativity at the quantum level, a quantum superposition must be broken and the system in question collapses into one of the possible states. In other words, Penrose proposes that a quantum state remains in superposition until the difference in spacetime curvature reaches a significant level. Penrose suggests that the wave function cannot sustain superposition beyond a certain energy difference between quantum states. When the energy difference exceeds a threshold, the wave function collapses to a single state based on its original amplitude. An interesting point is that observers do not play a special role in the collapse process. According to these statements, macroscopic systems cannot exist in multiple places simultaneously due to the large energy difference. Microscopic systems (such as electrons) can remain in superposition for longer (thousands of years) until their spacetime curvature separation reaches the collapse threshold. In Penrose’s approach too, an ensemble of spacetimes is inevitably considered. Since, the superposition of spacetimes is illegal and a well-defined timelike Killing vector can not be defined in this situation, the system must collapse into a particular state.
Penrose also mentions another interesting conceptual point. He argues that in order to establish the principle of equivalence at the quantum level, a unitary transformation should relate the wave functions of an accelerated observer (Newtonian observer) and an inertial observer (Einsteinian observer) together. The nonlinear dependence of the phase of this transformation on time, makes the concept of ”positive frequency” different for two observers. And the observers experience different spacetimes. The aforementioned transformation is as follows:
@@
Where, $\psi(\mathbf{x}^{\prime},t)$ and $\phi(\mathbf{x},t)$ are the wave functions of the Newtonian and Einsteinian observers respectively. This is the same physical concept that we are familiar with in the theory of quantum fields in curved spacetime, during which the accelerated observer does not see the quantum field as a pure field, but sees it as a mixture of states, i.e. a statistical system with a particular temperature. This particular temperature is called the Unruh temperature. During the wave function reduction, a pure state transforms to a mixture of states too. Therefore, a concept similar to the Unruh temperature may also be derived from this phenomenon. These concepts are also discussed in Refs. [16, 17]. In Ref. [18], we have shown that it is possible to define a temperature which has a close resemblance to the Unruh temperature in the problem of wave function reduction. In order to study the gravitational reduction of the wave function in the framework of standard quantum mechanics, Refs.[19, 20, 21, 22, 23, 24] are very suitable.
Here, it is appropriate to mention the general features of the Bohmian quantum mechanics. Bohmian quantum mechanics is an approach in which the particle has a deterministic quantum dynamics, and it is possible to consider trajectories for the particles. In the non-relativistic Bohmian quantum mechanics, a particle moves along a well-defined trajectory in the ensemble with the velocity
@@
where $\mathbf{p}=\nabla S(x,t)$ is the momentum of the particle. The meaning of $\mathbf{x}=\mathbf{x}(t)$, is that the particle moves along a trajectory of the ensemble which is described by $\mathbf{x}(t)$. Here, $S(x,t)$ is the principal function of the particle. Newton’s second law for the quantum motion of the particle in an external potential, $U$, takes the form
@@
Here, $\mathcal{R}$, denotes the amplitude of the wave function which is represented in the polar form
@@
The quantum potential has non-local property and is responsible for the quantum behavior of the particle[25, 26, 27, 28, 29, 30]. For example, the dispersion of a wave packet can be related to the quantum force, which is the derivative of the quantum potential. In the famous two-slit experiment, it can be shown that the trajectories of particles passing through two slits are affected by the quantum force in such a way that the place where the particles collide on the screen is observed in the form of interference fringes.
Despite the approaches that exist in the standard quantum mechanics to gravity-induced wave function reduction, we intend to investigate this issue in the framework of Bohmian quantum mechanics. Because in the Bohmian quantum mechanics, it is possible to define the quantum motion for the particle and related dynamical concepts can be used in a more obvious way. Therefore, it is possible to visualize the problem more intuitively. For example, in the investigation of the gravitational reduction of the wave function in the standard quantum mechanics, an ensemble of spacetimes is introduced. But, in the Bohmian quantum mechanics, an ensemble of trajectories is considered only in a single background. We will see how the concepts of trajectory and force will help us to easily obtain the critical quantities in the problem. The reduction time of the wave function and its critical width will be obtained using the quantum dynamics of the particle or object. In previous studies, we considered the matter for a point particle[31, 32, 33]. But, in this paper, we will generalize it to an object. Because what passes from the quantum world to the classical world is no longer a fundamental particle. So we felt the need to generalize our previous studies. In addition, we will further investigate the issue and obtain more accurate results.
@@ 2 The quantum motion and transition condition
In this section we provide a geometric approach for gravity-induced wave function reduction using Bohmian trajectories. We have done it before for a particle. Now let’s take a quick look at this approach and use the results for an object. It can be shown that the Bohmian trajectories do not cross each other[28]. Thus, we can consider an ensemble of non-crossing trajectories in configuration space where an arbitrary trajectory $\gamma$ is parametrized by the parameter $t$ and the particle can choose one of these trajectories and move along it with the velocity vector $\mathbf{u}^{\alpha}$.
@@
where, $U_{\text{cl}}$ denotes the classical potential energy of the particle or object. The quantum gravitational energy is the gravitational energy of the quantum distribution of the particle or object in configuration space.
Along the curves parametrized by $s$, we use $\frac{\partial}{\partial s}=\xi^{\beta}\frac{\partial}{\partial x^{\beta}}$ which helps us to write, relation (13) in the form,
@@
The terms in the parenthesis are quantum acceleration and quantum gravitational acceleration, respectively. It is obvious that to have zero deviation acceleration or parallel trajectories, we must have
@@
For a gravity-dominant regime in which the quantum gravitational force overcomes the quantum force, we expect the trajectories to be closer together, and in a quantum-dominant regime in which the quantum force overcomes the quantum gravitational force, the trajectories of the ensemble diverge[33]. Relation (17) will be used to obtain a criterion for the transition from the quantum world to the classical world. The condition $\nabla Q=-\nabla U_{\text{QG}}$, defines the transition regime in this context. We shall see that the average of the relation (17) gives the critical width of the wave packet for the transition between the quantum and classical domains.
3 The transition condition for a point particle
First, we take a look at the dynamics of a point particle. Then, we shall generalize it to include an object. We want to investigate the issue of under what conditions an object is macroscopic or microscopic. A macroscopic object is governed by Newton’s laws at the non-relativistic limit. While a microscopic object obeys the Schrödinger equation and exhibits quantum behavior. We first examine this for a point particle. When the quantum force that tends to diverge the particle’s position is balanced with the gravitational force that tends to converge the particle’s position, we expect to find a condition for the transition from the quantum world to the classical world. We conclude this result from the equation (15). The interesting thing is that when we reach the boundary of the classical and quantum worlds, both gravitational and quantum forces are present and there is no sharp distinction between the classical and quantum worlds. In other words, two gravitational and quantum forces are always present. But, for macroscopic objects, gravity dominates and for microscopic objects, the quantum force dominates.
We assume the particle is guided by a Gaussian wave packet. Here, we mention two points. First, we consider a homogeneous gravitational field in which the form of the wave packet remains Gaussian. In fact, it can be shown that in a short-time estimation, the form of the Gaussian wave packet remains unchanged with a good approximation[28, 33]. In other words, consider a free wave packet in a frame with coordinates $(\mathbf{x},t)$ and another frame with coordinates $(\mathbf{x}^{\prime},t)$ in which a homogeneous gravitational field $\mathbf{g}$, is present. It can be shown that the form of a Gaussian wave packet under the coordinate transformation $\mathbf{x}^{\prime}=\mathbf{x}+\frac{1}{2}\mathbf{g}t^{2}$ is invariant in a homogeneous gravitational field[33]. We shall see that the gravitational force and potentials in our study depend on the width of the wave packet and, for the short times, the width of the wave packet can be considered constant. The second point is that we perform calculations in spherical coordinates for the sake of simplicity. It is not difficult to show that the amplitude of a Gaussian wave packet in spherical coordinates takes the following form,

### v2

@@ Abstract
We present a novel perspective on gravity-induced wave function reduction using Bohmian trajectories. This study examines the quantum motion of both point particles and objects, identifying critical parameters for the transition from quantum to classical regimes. By analyzing the system’s dynamics, we define the reduction time of the wave function through Bohmian trajectories, introducing a fresh viewpoint in this field. Our findings align with results obtained in standard quantum mechanics, confirming the validity of this approach.
@@ 1 Introduction
The reduction of the wave function, a long-standing unresolved issue in quantum physics, has been studied through various approaches. For instance, GRW theory suggests that wave function collapse occurs spontaneously and randomly over time [1, 2]. Decoherence theory explains how interactions with the environment cause a quantum system to lose coherence, thereby behaving in a classical manner [3, 4, 5]. The many-worlds interpretation posits that all possible outcomes of a measurement actually occur, each in its own separate branch of the universe, resulting in a vast multiverse where every possible result is realized in some branch [6, 7]. In our study, we explore the gravitational reduction of the wave function. This approach has been extensively examined within the context of standard quantum mechanics (see Ref.[8]). However, we aim to investigate this topic through the lens of Bohmian quantum mechanics, which offers greater intuition and insight into the underlying processes.
In gravity-induced wave function reduction, the mass of an object serves as a criterion for determining the boundary between the classical and quantum worlds. In this interpretation, the self-gravity of the object or the curvature of spacetime caused by its mass directs the object to a specific state. This concept, first formulated around sixty years ago [9, 10, 11], remains relevant. Macroscopic objects adhere to the laws of classical mechanics, while the behavior of microscopic objects is governed by the Schrödinger equation. Thus, defining a critical mass is essential to demarcate the boundary between the classical and quantum realms.
The concept of self-gravity should not be confusing. Consider a particle or object to which a wave packet is attributed. According to the principles of quantum mechanics, this particle or object can be in multiple places simultaneously, meaning we have a superposition of different states of the particle. Within this probability space, a distribution of matter locations can be considered with self-gravitational energy. Since this distribution is rooted in the uncertainty principle, we refer to its associated energy as ’quantum gravitational energy.’ The quantum superposition is unstable in the presence of such a gravitational field and is disrupted when the mass of the particle or object reaches the critical mass. When the superposition breaks down, the particle or object moves towards greater certainty. The timescale between breaking the superposition and the system settling into a specific state is often referred to as the reduction time of the wave function. This concept aligns closely with the ideas of Penrose and Karolyhazi, while the damping time proposed by Diosi is also relevant to this concept. This topic has been extensively studied in standard quantum mechanics. Now, we will examine it within the framework of Bohmian quantum mechanics, where definitions such as trajectory, force, and velocity are meaningful. Our goal is to provide a fresh perspective on the concept of reduction time. Despite numerous achievements in standard quantum mechanics, an ontological and deterministic description of the problem remains elusive. We cannot predict with certainty which of the possible states the quantum system will occupy after measurement or the breaking of superposition by gravity, and we rely solely on statistical predictions. This unpredictability persists in Bohmian quantum mechanics as well.
@@
First, let’s explore the gravitational approaches that exist in standard quantum mechanics. The first significant work in this area is by Karolyhazi, who considered an uncertainty for the metric of spacetime. To achieve the classical regime, the uncertainty in spacetime is minimized in the presence of a scalar field. This leads to the critical conditions necessary for the transition from the quantum domain to the classical domain. Karolyhazy’s relations for the critical width and reduction time of an elementary particle are given by
@@
Here, the reduction time is the time it takes for a quantum state to undergo spontaneous localization, leading to a classical-like behavior. For a proton, one can find $\sigma_{c}\approx 10^{22}m$ and $\tau\approx 10^{53}s$. For an object of mass $M$ and size $R$, Karolyhazy derives the following relation:
@@
This relation helps in determining the critical conditions necessary for the transition from the quantum domain to the classical domain based on the object’s mass and size. For a tennis ball of radius $\sim 4cm$ and mass $\sim 57g$, the critical width and reduction time are approximately $\sigma_{c}\sim 10^{-17}cm$ and $\tau\sim 10^{-6}s$ respectively. For details, see[9, 10, 11].
The gravitational approach of Diosi also determines a critical width for the transition from the quantum to the classical world. Diosi’s first significant work is an investigation through the Schrödinger-Newton equation, a nonlinear equation with a self-gravitational potential. It is believed that such a nonlinear term breaks down the unitary transformation and transforms a pure state into a mixture of states. By using the Schrödinger-Newton equation, the critical width for the transition is obtained. Subsequent works by Diosi deal with the concept of intrinsic quantum imprecision in the spacetime metric. By studying the fluctuations of the gravitational field when measured by a quantum probe over a period of time, the damping time or reduction time of a quantum state in the presence of gravity is defined [12, 13, 14]. The critical widths for the transition from the quantum world to the classical world of an object in Diosi’s approach are as follows:
@@
In the above relation, a macro object refers to a macroscopic body that exhibits classical behavior and does not exhibit significant quantum superposition over macroscopic distances. A micro object exhibits quantum behaviors such as superposition and entanglement. We expect that a micro object to be in different places or states at the same time and the width of its wave packet to be greater than its classical size. For a proton, the critical width is about $10^{4}m$, which is really greater than its classical size,.i.e $10^{-15}m$. The damping time or the reduction time for a proton is about $10^{15}s\approx 32000000$ years, which indicates that we can not see the collapse of the proton normally.
The approach of Penrose is very fundamental. In Ref. [15], an effort is made to bring quantum mechanics closer to the principles of general relativity. Penrose shows that in order to establish the principles of general relativity at the quantum level, a quantum superposition must be broken and the system in question must collapse into one of the possible states. In other words, Penrose proposes that a quantum state remains in superposition until the difference in spacetime curvature reaches a significant level. Penrose suggests that the wave function cannot sustain superposition beyond a certain energy difference between quantum states. When the energy difference exceeds a threshold, the wave function collapses to a single state based on its original amplitude. An interesting point is that observers do not play a special role in the collapse process. According to these statements, macroscopic systems cannot exist in multiple places simultaneously due to the large energy difference. Microscopic systems (such as electrons) can remain in superposition for longer (thousands of years) until their spacetime curvature separation reaches the collapse threshold. In Penrose’s approach, an ensemble of spacetimes is inevitably considered. Since the superposition of spacetimes is illegal and a well-defined timelike Killing vector cannot be defined in this situation, the system must collapse into a particular state.
Penrose highlights another intriguing conceptual point. He argues that to establish the principle of equivalence at the quantum level, a unitary transformation must relate the wave functions of an accelerated observer (Newtonian observer) and an inertial observer (Einsteinian observer). The nonlinear dependence of the phase of this transformation on time causes the concept of ’positive frequency’ to differ for the two observers, resulting in them experiencing different spacetimes. The aforementioned unitary transformation is as follows:
@@
Where, $\psi(\mathbf{x}^{\prime},t)$ and $\phi(\mathbf{x},t)$ are the wave functions of the Newtonian and Einsteinian observers respectively. This is the same physical concept we are familiar with in the theory of quantum fields in curved spacetime, during which an accelerated observer does not see the quantum field as a pure field but as a mixture of states, i.e., a statistical system with a particular temperature. This particular temperature is called the Unruh temperature. During wave function reduction, a pure state also transforms into a mixture of states. Therefore, a concept similar to the Unruh temperature may also be derived from this phenomenon. These concepts are discussed in Refs. [16, 17]. In Ref. [18], we have shown that it is possible to define a temperature that closely resembles the Unruh temperature in the problem of wave function reduction. To study the gravitational reduction of the wave function within the framework of standard quantum mechanics, Refs. [19, 20, 21, 22, 23, 24] are very suitable.
Here, it is appropriate to mention the general features of Bohmian quantum mechanics. Bohmian quantum mechanics is an approach in which the particle has a deterministic quantum dynamics, allowing us to consider trajectories for the particles. In non-relativistic Bohmian quantum mechanics, a particle moves along a well-defined trajectory in the ensemble with the velocity
@@
where $\mathbf{p}=\nabla S(x,t)$ is the momentum of the particle. The meaning of $\mathbf{x}=\mathbf{x}(t)$ is that the particle moves along a trajectory of the ensemble which is described by $\mathbf{x}(t)$. Here, $S(x,t)$ is the principal function of the particle. Newton’s second law for the quantum motion of the particle in an external potential, $U$, takes the form
@@
where $\mathcal{R}$ is the amplitude of the wave function which can be represented in the following polar form
@@
where $U$ is the classical potential energy. The quantum potential has non-local property and is responsible for the quantum behavior of the particle[25, 26, 27, 28, 29, 30]. For example, the dispersion of a wave packet can be related to the quantum force, which is the derivative of the quantum potential. In the famous two-slit experiment, it can be shown that the trajectories of particles passing through two slits are affected by the quantum force in such a way that the places where the particles collide on the screen form interference fringes.
In Bohmian quantum mechanics, it is possible to study the quantum motion of the particle or object, which allows for a more intuitive visualization of the problem. We will see how concepts like trajectory and force help us easily obtain the critical quantities for the transition between the quantum and classical worlds. The reduction time of the wave function and its critical width will be obtained using the quantum dynamics of the particle or object. In previous studies, we considered the matter for a point particle [31, 32, 33]. However, in this paper, we will generalize it to an object, as what transitions from the quantum world to the classical world is no longer an elementary particle. In the framework of Bohmian quantum mechanics, there is no sharp distinction between the quantum and classical worlds. Instead, there is a relative interplay between the quantum force and gravity. The gravitational force characterizes the classical world, while the quantum force is intrinsic to the quantum world, with its origin still not fully understood. When the gravitational force dominates, the behavior of the particle or object exhibits less uncertainty. Conversely, when the quantum force is dominant, uncertainty becomes more prominent. In other words, the quantum force is not exactly zero in the classical world, just as the gravitational force is not zero in the quantum world.
@@ 2 The Quantum Motion and Transition Condition
In this section we provide a geometric approach for gravity-induced wave function reduction using Bohmian trajectories. We have done this before for a point particle. Now let’s take a quick look at this approach and use the results for an object. It can be shown that the Bohmian trajectories do not cross each other[28]. Thus, we can consider an ensemble of non-crossing trajectories in configuration space where an arbitrary trajectory $\gamma$ is parametrized by the parameter $t$ and the particle can choose one of these trajectories and move along it with the velocity vector $\mathbf{u}^{\alpha}$. See Fig.1.
@@
Here, $U_{\text{cl}}$ denotes the classical potential energy of the particle or object. As mentioned in the introduction, the quantum gravitational energy refers to the gravitational energy of the quantum distribution of the particle or object in configuration space.
Along the curves parametrized by $s$, we use $\frac{\partial}{\partial s}=\xi^{\beta}\frac{\partial}{\partial x^{\beta}}$ which helps us to write relation (13) in the form:
@@
The terms in the parenthesis are quantum acceleration and quantum gravitational acceleration, respectively. In a gravity-dominant regime, where the quantum gravitational force overcomes the quantum force, we expect the trajectories to be closer together. Conversely, in a quantum-dominant regime, where the quantum force overcomes the quantum gravitational force, the trajectories of the ensemble diverge [33]. It is obvious that to have zero deviation acceleration or parallel trajectories, we must have
@@
This relation will be used to obtain a criterion for the transition from the quantum world to the classical world. The condition $\nabla Q=-\nabla U_{\text{QG}}$ defines the transition regime in this context. We shall see that the average of the relation (17) gives the critical width of the wave packet for the transition between the quantum and classical domains.
2.1 The Transition Condition for a Point Particle
Through the relation (15), we obtain the condition (17), which states that when the quantum force and the quantum gravitational forces are balanced, any two nearby trajectories in the ensemble remain parallel. In this situation, we expect that the associated wave packet does not spread and remains a stationary wave packet. Thus, the average of relation (17) may lead us to the critical values for transition.
We assume the particle is guided by a Gaussian wave packet. If we consider a homogeneous gravitational field, the form of the wave packet remains Gaussian. This is possible when we consider the problem in a short-time estimation [28, 33]. We utilize the spherical form of the amplitude of the Gaussian wave packet for the sake of simplicity, which has the following form: